# PROTOCOL FOR AN OBSERVATIONAL STUDY ON THE EFFECTS OF COMBINATIONS OF ADVERSE CHILDHOOD EXPERIENCES ON ADULT DEPRESSION

### v1

@@ 1. Introduction
Adverse childhood experiences (ACEs) describe specific categories of early life hardship or harm that occur before an individual reaches adulthood. Although these can take various forms, they are typically categorized into ten types involving abuse (physical, emotional, or sexual) and neglect (physical or emotional), as well as household challenges including mental illness, substance abuse, intimate partner violence, parental absence due to divorce or separation, or incarceration. Initially investigated in a landmark study conducted by the Centers for Disease Control and Prevention (CDC) and Kaiser Permanente, ACEs are highly prevalent and exhibit a strong dose-response relationship with numerous negative health and social outcomes in adulthood (Felitti et al. [1998]; Dube et al. [2003]; Anda et al. [2006]). Exposure to individual ACEs has been repeatedly associated with a range of adverse physical and emotional health consequences later in life. One study indicated that childhood emotional abuse increases the risk of lifetime depressive disorders, with adjusted odds ratios (ORs) of 2.7 for women and 2.5 for men (Chapman et al. [2004]). Depression also appears to be closely related to having experienced parental alcohol abuse (Anda et al. [2002]). Moreover, physical and verbal abuse have been found to be strongly associated with body weight issues and obesity (Williamson et al. [2002]).
@@
Unfortunately, when children are exposed to one type of adversity, there is a high likelihood of exposure to additional ACEs (Dong et al. [2004]; Felitti [2002]). Numerous large-scale investigations have confirmed that these detrimental experiences can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al. [1998]; Chapman et al. [2004]). Accordingly, previous literature has examined closely the number of categories of ACEs, often referred to as ACE score. For instance, one study using data from 14 states revealed substantial variation across states in the percentage change in health outcomes for those exposed to at least four ACEs compared to those with fewer than four ACEs (Waehrer et al. [2020]). Physically, individuals with four or more ACEs face elevated risks of heart disease, liver disease, kidney disease, asthma, Alzheimer’s disease, stroke, metabolic disorders, obesity, diabetes, and certain cancers (for Disease Control and Prevention [2017]; for Youth Wellness [2014]; Merrick [2019]; Hughes et al. [2017]; Waehrer et al. [2020]). On the emotional and psychological side, ACE score strongly correlates with mental health issues such as depression, anxiety, and suicidal ideation (Hillis et al. [2004]; Chapman et al. [2004]). Meta-analyses further reinforce that an accumulation of at least 4 ACEs correlates with higher risks of chronic disease, poor mental health, and negative social outcomes, demonstrating the need for strategies to prevent ACEs and lessen their long-term repercussions (Hughes et al. [2017]).
A recent study challenged the overall ACE score and advised to break down the measure back into specific items (Fitzgerald and Bishop [2024]). Gaps also remain in our understanding of how the accumulation of ACEs influences health. We are therefore interested in which categories of ACEs—and which specific combinations of ACEs (e.g., physical abuse + sexual abuse, or physical abuse + emotional neglect + substance abuse)—are most harmful and warrant primary care. We focus our analysis on depression, one of the most prevalent and burdensome mental health conditions globally. Consequently, our focus is to evaluate the effect of different combinations of specific ACE items on depression and identifying a higher-risk group of sets of ACEs.
@@
We will examine the impacts of combinations of ACEs on adult depression using data from the Behavioral Risk Factor Surveillance System (BRFSS) 2023, an ongoing health-related telephone survey system administered by the CDC with a random sample of over 400,000 individuals across the United States in 2023. This data includes detailed information from ten states that administered the CDC’s Adverse Childhood Experiences module in 2023. We can obtain all ten ACE categories proposed in the CDC-Kaiser study (Felitti et al. [1998]), and there are over 50,000 individuals with complete records of all ACE items. Among them, around 17,000 reported zero ACEs while the rest were exposed to at least one ACE. This large sample size will help facilitate powerful data analysis and explore the influencing mechanism of different ACEs combinations. The BRFSS also collected measurements of variables that may confound the relationship between exposure to ACEs and depression, such as those related to personal demographics. We will utilize those variables to form separate strata for treated and control individuals, enabling us to compare the outcomes of similar individuals in hopes of ameliorating confounding biases.
@@
We employ a novel statistical design called data turnover (Bekerman et al. in preparation, see Bekerman et al. [2024a] for a description) to conduct this analysis. By integrating both qualitative and quantitative insights from data exploration, data turnover enables a single team of statisticians and domain experts to evaluate evidence from multiple data splits. Exploratory data analysis (EDA) allows us to inspect the data before making potentially questionable assumptions, detect anomalies or outliers that do not represent natural variation in the data, and formulate new hypotheses. It also helps identify variables that may not measure what we initially intended, indicating the need for a better measure or refined, data-informed hypotheses (Bekerman et al. [2024b]). Beyond exploration, replication is critical in observational studies for reinforcing the credibility of findings and sustaining trust in scientific knowledge. Previous work has combined some—but not all—aspects of exploration and replication within the same observational study. For instance, some methods allow for exploration without replication (e.g., Cox [1975]; Heller et al. [2009]), whereas others enable replication without exploration (e.g., Zhao et al. [2018]; Karmakar et al. [2019]). There is also research that supports both exploration and replication but requires two independent investigative teams (Roy et al. [2022]). In contrast, data turnover makes it possible to conduct both exploration and replication using a single research group.
@@
In observational studies like ours, where treatment assignment is not under our direct control, obtaining consistent findings across multiple groups that may have different assignment mechanisms reinforces the evidence that the treatment itself is truly responsible for the observed effects (Rosenbaum [2001]; Rosenbaum [2015]). By analyzing data from two distinct subpopulations of BRFSS 2023 participants—residing in states that voted for the Democratic candidate in the 2024 US presidential election (”blue states”) and those residing in states that voted for the Republican candidate (”red states”) —we can assess the replicability of our findings. We split our data into blue states and red states, recognizing that these political orientations often shape distinct social, economic, and health care policies. A study focusing on population in East and West Germany has shown that socio-political context is a determinant of childhood maltreatment (Ulke et al. [2021]). In our context, for example, blue states may allocate more resources toward public health programs (e.g., adopt medicaid expansion) or have broader social safety nets, whereas red states may emphasize different strategies for healthcare funding and child protective services. Such policy and cultural variations can influence both the prevalence of ACEs and the mechanisms by which individuals are exposed to them, potentially resulting in different patterns of ACEs across these two groups. Demonstrating consistent results in both blue and red states thus enhances confidence that our findings are robust across varied socio-political contexts. By employing data turnover, we can predetermine how one subgroup will guide analysis of the other, then apply exploratory data analysis to the second subgroup to inform how we draw inferences about the first. Any outcome that shows a significant effect in both groups is defined as a “replicable” finding. (Bogomolov and Heller [2023]). In Section 5, we detail how we apply this non-random cross-screening method to detect replicable outcomes and test the global null hypothesis, i.e., identifying outcomes that are affected in at least one of the two subpopulations.
@@
The remainder of the document is organized as follows. We introduce the BRFSS dataset and describe our data processing in Section 2. In Section 3, we report our candidate sets of combinations. In Section 4, we outline our stratification methodology. Finally, we conclude with a description of our data analysis procedure and explanation of several specifics of implementation in Section 5.
@@ 2. The Behavioral Risk Factor Surveillance System (BRFSS)
This study utilizes data from BRFSS, an ongoing health-related telephone survey system administered by the CDC. BRFSS collects comprehensive data on a wide range of health-related behaviors, chronic health conditions, healthcare access, and preventive service usage from the non-institutionalized adult population (18 years and older). The survey employs a dual-frame sampling methodology, using both landline and cellular telephone surveys to ensure representativeness of the data. The dataset collected in 2023 is the latest data available, has 433,323 observations from 48 states, the District of Columbia, Guam, Puerto Rico, and the US Virgin Islands, and includes topics such as health status, chronic conditions, health behaviors, healthcare access, and preventive services, as well as optional modules including social determinants of health, cancer screenings, and cognitive decline.
@@
For the purpose of our study, we focus on the 10-state [^1] Delaware, Florida, Georgia, Missouri, Nevada, New Jersey, Oregon, Rhode Island, Tennessee, Virginia. sample that administered the CDC’s ACEs module in 2023. Following previous studies utilizing the BRFSS data (Bhan et al. [2014]; Gilbert et al. [2015]; Waehrer et al. [2020]), we collapse different questions into particular indicators. More details are provided in table 1. In total, we create 10 ACEs items from the BRFSS data (see table 2), which is consistent with the categorization in CDC-Kaiser study (Felitti et al. [1998]). It is worth noting that we first follow those criteria for data pre-processing, but these steps could be adjusted when we carry out exploratory data analysis in the data turnover..
Indicator Questions Collapsed Encoding Method
@@
Any Sexual Abuse Number of times forced to touch, be touched, or have sex with anyone at least five years older or an adult Collapsed into a single indicator for any sexual abuse
Household Substance Use Alcohol and illegal drug use in the household Combined into a single indicator for household substance use
@@
Physical Abuse Frequency of physical abuse Encoded as 1 if the answer is “once” or “more than once”
@@
Verbal Abuse Frequency of verbal abuse Encoded as 1 if the answer is “once” or “more than once”
Parental Divorce Response to parental divorce question Affirmative response encoded as 1
@@
Living with a Depressed or Mentally Ill Person Response to question about living with a mentally ill person Affirmative response encoded as 1
@@
Living with an Incarcerated Person Response to question about living with an incarcerated person Affirmative response encoded as 1
@@
Physical Violence Between Parents Frequency of physical violence between parents Affirmative response encoded as 1
@@
Safety Did an adult make you feel safe and protected ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
@@
Basic Needs Met Did an adult make sure basic needs were met ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
@@
After processing the data, we restrict our sample to those who do not have any missing values in any ten ACEs variables. We also omit those individuals who do not have the appropriate measurement for depression. For any missing family background, we directly use the imputation value provided by the BRFSS 2023. In total, there are 49,811 remaining individuals, among whom 23,506 individuals are from blue states [^2] Delaware, New Jersey, Oregon, Rhode Island, Virginia (as established in the 2024 United States presidential election) and 26,305 individuals from red states [^3] Florida, Georgia, Missouri, Nevada, Tennessee. In the part of data with individuals from blue states, 8,076 individuals do not have any ACEs while 15,512 were exposed to at least one ACE. In the part of data with individuals from blue states, 9,163 individuals do not have any ACEs while 17,231 were exposed to at least one ACE.
3. Pre-specified Candidate Sets of ACEs
We collect ten variables of ACEs from the BRFSS 2023 dataset. To examine the diverse effects of combinations of ACEs on depression, we first construct candidate sets of ACEs that we are interested in. Since previous literature has identified four ACEs as a critical threshold, beyond which the effects appear to intensify significantly, and that individuals with at least four ACEs are at an increased risk for chronic disease, poor mental health, and negative social outcomes (Hughes et al. [2017]; Waehrer et al. [2020]), we consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs categories. We thus have 385 possible combinations of ACEs formed from ten candidate ACEs categories (see Appendix A).
4. Stratification
A previous study (Waehrer et al. [2020]) used BRFSS 2009 & 2012 to examine whether the relationship between ACEs and health outcomes, such as smoking, heavy drinking, and obesity, was similar across states, and controlled for baseline covariates including race/ethnicity, sex, and age. They also controlled for education, marital status and employment, which are post-treatment variables and likely to be affected by ACEs. Cox [1958] refers to these variables as “concomitant observations” and writes: “The essential point in our assumptions about these observations is that the value for any unit must be unaffected by the particular assignment of treatments to units actually used. In practice this means that either (a) the concomitant observations are taken before the assignment of treatments to units is made; or (b) the concomitant observations are made after the assignment of treatments, but before the effect of treatments has had time to develop…; or (c) we can assume from our knowledge of the nature of the concomitant observations concerned, that they are unaffected by the treatment.” If adjustments are made for variables that are affected by the treatment, then adjustments may remove part or all of the effect of the treatment, and hence estimators that adjust for a concomitant variable that has been affected by the treatment are generally biased (Rosenbaum [1984]). Therefore, we do not adjust for education, employment and marital status in our study.
Given the relatively small number of covariates, we seek to mitigate potential confounding bias by stratifying participants on their recorded baseline characteristics (race/ethnicity, age, sex) and then comparing those exposed to specific ACE combinations (the “treated” group) with those reporting no ACE exposure (the “control” group) within each stratum. Each hypothesis contrasts a particular ACE combination against having no ACEs. Individuals who have any additional ACEs beyond the combination of interest remain in the treated group. For example, if our focus is on the effect of substance abuse, participants exposed to both substance abuse and physical abuse are still classified under “substance abuse.” We conduct these stratifications separately for individuals residing in blue states and those in red states, resulting in $385\times 2$ distinct stratified analyses.
Specifically, we stratify participants by race/ethnicity, including White (non-Hispanic), Black (non-Hispanic), Hispanic, and Other (non-Hispanic); by age, with groups spanning 18-24, 25-34, 35-44, 45-54, 55-64, and 65+; and by sex, distinguishing between male and female participants. This results in a total of 48 distinct strata for both the blue and red subgroups. In the blue subgroup, the stratum with the lowest sample size contains 61 participants, with a median of 171. In the red subgroup, the stratum with the lowest sample size has 60 participants, with a median of 161. Detailed sample sizes for each stratum of both subgroups are provided in the appendix B.

### v2

@@ 1. Introduction
Adverse childhood experiences (ACEs) describe specific categories of early life hardship or harm that occur before an individual reaches adulthood. Although these can take various forms, they are typically categorized into ten types involving abuse (physical, emotional, or sexual) and neglect (physical or emotional), as well as household challenges including mental illness, substance abuse, intimate partner violence, parental absence due to divorce or separation, or incarceration. ACEs were initially investigated in a landmark study conducted by the Centers for Disease Control and Prevention (CDC) and Kaiser Permanente, and are highly prevalent and exhibit a strong dose-response relationship with numerous negative health and social outcomes in adulthood (felitti1998relationship; dube2003impact; anda2006enduring).
Exposure to individual ACEs has been repeatedly associated with a range of adverse physical and emotional health consequences later in life. One study has found that childhood emotional abuse increases the risk of lifetime depressive disorders, with adjusted odds ratios (ORs) of 2.7 for women and 2.5 for men (chapman2004adverse). Depression also appears to be closely related to having experienced parental alcohol abuse (anda2002adverse). Moreover, physical and verbal abuse have been found to be strongly associated with body weight issues and obesity (williamson2002body).
@@
Unfortunately, when children are exposed to one type of adversity, there is a high likelihood of exposure to additional ACEs (dong2004interrelatedness; felitti2002relationship). Numerous large-scale investigations have confirmed that these detrimental experiences can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (felitti1998relationship; chapman2004adverse). Accordingly, previous literature has closely examined the number of categories of ACEs, often referred to as the ACE score. For instance, one study using data from 14 states revealed substantial variation across states in the percentage change in health outcomes for those exposed to at least four ACEs compared to those with fewer than four ACEs (waehrer2020disease). Physically, individuals with four or more ACEs face elevated risks of heart disease, liver disease, kidney disease, asthma, Alzheimer’s disease, stroke, metabolic disorders, obesity, diabetes, and certain cancers (CDC2017; CYW2014; merrick2019vital; hughes2017effect; waehrer2020disease). On the emotional and psychological side, the ACE score strongly correlates with mental health issues such as depression, anxiety, and suicidal ideation (hillis2004association; chapman2004adverse). Meta-analyses further reinforce that an accumulation of at least 4 ACEs correlates with higher risks of chronic disease, poor mental health, and negative social outcomes, demonstrating the need for strategies to prevent ACEs and lessen their long-term repercussions (hughes2017effect).
A growing literature cautions that the additive ACE score is a blunt tool for individual risk, as it treats all events as equal. Then, a child enduring repeated sexual abuse is assumed the same as exposure to parental divorce, which appears problematic on its face. The ACE score can encourage over-simplified claims about risk or causality, deterministic framing, and stigma (lacey2020practitioner). Methodologically, large cohort analyses show that an ACE score cutoff such as $\geq 4$ poorly distinguishes which specific individuals will develop later health problems (baldwin2021population; meehan2022poor). A cohort study of women in correctional custody showed that maltreatment alone is significantly associated with adult mental health while household dysfunction is not. Therefore, the authors challenged the overall ACE score and recommended decomposing the measure back into specific items (fitzgerald2024challenging). Recent pediatric guidance finds insufficient evidence to support routine, threshold-based ACE screening in clinical settings and supports trauma-informed, comprehensive assessment instead (austin2024screening). Gaps also remain in our understanding of how the accumulation of ACEs influences health. We are therefore interested in which categories of ACEs—and which specific combinations of ACEs (e.g., physical abuse + sexual abuse, or physical abuse + emotional neglect + substance abuse)—are most harmful and warrant particular attention in primary care settings. We focus our analysis on depression, one of the most prevalent and burdensome mental health conditions globally. Consequently, our focus is to evaluate the effect of different combinations of specific ACE items on depression and to identify a higher-risk group of sets of ACEs.
@@
We will examine the impacts of combinations of ACEs on adult depression using data from the Behavioral Risk Factor Surveillance System (BRFSS) 2023, an ongoing health-related telephone survey system administered by the CDC with a random sample of over 400,000 individuals across the United States in 2023. These data include detailed information from ten states that administered the CDC’s Adverse Childhood Experiences module in 2023. We observe all ten ACE categories proposed in the CDC-Kaiser study (felitti1998relationship), and there are over 50,000 individuals with complete records of all ACE items. Among them, around 17,000 reported zero ACEs while the rest were exposed to at least one ACE. This large sample size will help support powerful data analysis and facilitate the exploration of influencing mechanism of different ACEs combinations.
@@
We employ a novel statistical design called data turnover (Bekerman et al. in preparation, see bekerman2024protocol for a description) to conduct this analysis. By integrating both qualitative and quantitative insights from data exploration, data turnover enables a single team of statisticians and domain experts to evaluate evidence from multiple data splits. Exploratory data analysis (EDA) allows us to inspect the data before making potentially questionable assumptions, detect anomalies or outliers that do not represent natural variation in the data, and formulate new hypotheses. It also helps identify variables that may not measure what we initially intended, indicating the need for a better measure or refined, data-informed hypotheses (bekerman2024planning). Beyond exploration, replication is critical in observational studies for reinforcing the credibility of findings and sustaining trust in scientific knowledge. Previous work has combined some—but not all—aspects of exploration and replication within the same observational study. For instance, some methods allow for exploration without replication (e.g., cox1975note; heller2009split), whereas others enable replication without exploration (e.g., zhao2018cross; karmakar2019integrating). There is also research that supports both exploration and replication but requires two independent investigative teams (roy2022protocol). In contrast, data turnover makes it possible to conduct both exploration and replication using a single research group.
@@
In observational studies like ours, where treatment assignment is not under our direct control, obtaining consistent findings across multiple groups that may have different assignment mechanisms reinforces the evidence that the treatment itself is truly responsible for the observed effects (rosenbaum2001replicating; rosenbaum2015see). By analyzing data from two distinct subpopulations of BRFSS 2023 participants—residing in states that voted for the Democratic candidate in the 2024 US presidential election (”blue states”) and those residing in states that voted for the Republican candidate (”red states”) —we can assess the replicability of our findings. We split our data into blue states and red states, recognizing that these political orientations often shape distinct social, economic, and health care policies. A study focusing on the population in East and West Germany has shown that socio-political context is a determinant of childhood maltreatment (ulke2021socio). In our context, for example, blue states may allocate more resources toward public health programs (e.g., adopt Medicaid expansion) or have broader social safety nets, whereas red states may emphasize different strategies for healthcare funding and child protective services. Such policy and cultural variations can influence both the prevalence of ACEs and the mechanisms by which individuals are exposed to them, potentially resulting in different patterns of ACEs across these two groups. Demonstrating consistent results in both blue and red states thus enhances confidence that our findings are robust across varied socio-political contexts. By employing data turnover, we can predetermine how one subgroup will guide analysis of the other, then apply exploratory data analysis to the second subgroup to inform how we draw inferences about the first. Any outcome that shows a significant effect in both groups is defined as a “replicable” finding (bogomolov2023replicability). In Section 4, we detail how we apply this non-random cross-screening method to detect replicable outcomes and test the global null hypothesis, i.e., identifying outcomes that are affected in at least one of the two subpopulations.
@@
The remainder of the document is organized as follows. We introduce the BRFSS dataset and describe our data processing in Section 2. In Section 3, we outline our methodology. Finally, we conclude with a description of our data analysis procedure and an explanation of several specifics of implementation in Section 4.
@@ 2. The Behavioral Risk Factor Surveillance System (BRFSS)
This study uses data from BRFSS, an ongoing health-related telephone survey system administered by the CDC. BRFSS collects comprehensive data on a wide range of health-related behaviors, chronic health conditions, healthcare access, and preventive service usage from the non-institutionalized adult population (18 years and older). The survey employs a dual-frame sampling methodology, using both landline and cellular telephone surveys to ensure representativeness of the data. The dataset collected in 2023 is the latest data available, has 433,323 observations from 48 states, the District of Columbia, Guam, Puerto Rico, and the US Virgin Islands, and includes topics such as health status, chronic conditions, health behaviors, healthcare access, and preventive services, as well as optional modules including social determinants of health, cancer screenings, and cognitive decline.
@@
For the purpose of our study, we focus on the 10-state [^1] Delaware, Florida, Georgia, Missouri, Nevada, New Jersey, Oregon, Rhode Island, Tennessee, Virginia. sample that administered the CDC’s ACEs module in 2023. Following previous studies utilizing the BRFSS data (bhan2014childhood; gilbert2015childhood; waehrer2020disease), we collapse different questions into binary indicators of each ACE. More details are provided in Table 1. In total, we create 10 binary ACE items from the BRFSS data (see Table 2), which is consistent with the categorization in the CDC-Kaiser study (felitti1998relationship). Depression is coded Y=1Y{=}1 if the answer to ’(Ever told) (you had) a depressive disorder (including depression, major depression, dysthymia, or minor depression)?’ is ’Yes’ and 0 otherwise.
ACEs Item Questions Collapsed Encoding Method
@@
Any Sexual Abuse Number of times forced to touch, be touched, or have sex with anyone at least five years older or an adult Constructed by summing responses across multiple questions: all “never” encoded as 0, otherwise encoded as 1
Household Substance Use Alcohol and illegal drug use in the household Constructed by summing responses across multiple questions: all “never” encoded as 0, otherwise encoded as 1
@@
Physical Abuse Frequency of physical abuse ‘Once’ or ‘More than once’ encoded as 1, ‘None’ encoded as 0
@@
Verbal Abuse Frequency of verbal abuse ‘Once’ or ‘More than once’ encoded as 1, ‘None’ encoded as 0
Parental Divorce Response to parental divorce question ’Yes’ encoded as 1, ’No’ encoded as 0
@@
Living with a Depressed or Mentally Ill Person Response to question about living with a mentally ill person ’Yes’ encoded as 1, ’No’ encoded as 0
@@
Living with an Incarcerated Person Response to question about living with an incarcerated person ’Yes’ encoded as 1, ’No’ encoded as 0
@@
Physical Violence Between Parents Frequency of physical violence between parents ‘Once’ or ‘More than once’ encoded as 1, ‘None’ encoded as 0
@@
Emotional Neglect For how much of your childhood was there an adult in your household who made you feel safe and protected ‘Never’ and ‘A little of the time’ encoded as 1, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 0
@@
Physical Neglect For how much of your childhood was there an adult in your household who tried hard to make sure your basic needs were met ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
@@
After processing the data, we restrict our sample to those who do not have any missing values in any of the ten ACE variables. We also omit those individuals who do not have the appropriate measurement for depression. In total, there are 49,547 remaining individuals, among whom 23,390 individuals are from blue states [^2] Delaware, New Jersey, Oregon, Rhode Island, Virginia (as established in the 2024 United States presidential election) and 26,157 individuals from red states [^3] Florida, Georgia, Missouri, Nevada, Tennessee. In the part of data with individuals from blue states, 7,363 individuals do not have any ACEs while 16,027 were exposed to at least one ACE. In the part of data with individuals from red states, 8,479 individuals do not have any ACEs while 17,678 were exposed to at least one ACE.
3. Isotonic Subgroup Selection
We plan to use isotonic subgroup selection (ISS) to identify the high-risk group in our study (muller2025isotonic). Let $(X_{i},Y_{i})$, for $i=1,\dots,n$, be independent observations, where $X_{i}\in\mathbb{R}^{d}$ denotes covariates and $Y_{i}$ is the outcome of interest. In our setting, $Y_{i}$ is a binary indicator for depression, and $X_{i}$ consists of different combinations of ten ACE variables, each coded as binary. We assume a regression function $\eta(x)=\mathbb{E}(Y|X=x)$ and impose the coordinate-wise monotonicity assumption: $x\preceq x^{\prime}\implies\eta(x)\leq\eta(x^{\prime})$, where $x\preceq x^{\prime}$ means $x_{j}\leq x^{\prime}_{j}$ for all coordinates $j=1,\dots,d$. In practice, this monotonicity captures the expectation that exposure to additional ACEs should not reduce the probability of depression. For example, an individual reporting both physical abuse and parental substance use is assumed to have risk at least as high as someone reporting only parental substance use.
Fix a threshold $\tau$. Our target is the superlevel region $X_{\tau}(\eta)=\{\,x:\eta(x)\geq\tau\,\}$, which we interpret as the subgroup of covariate profiles whose expected probability of depression exceeds $\tau$. The goal of Isotonic Subgroup Selection (ISS) is to estimate a subset $\hat{A}$ that approximates $X_{\tau}(\eta)$ while controlling error rates in a principled way.
The ISS procedure consists of three main steps:
  1. (1)
    Hypothesis construction.
    For each $X_{i}$, we test $H_{0,i}:\eta(X_{i})<\tau\quad\text{versus}\quad H_{1,i}:\eta(X_{i})\geq\tau.$ Because $Y$ is binary, classification-based $p$-values are constructed using anytime-valid martingale methods for bounded outcomes. Intuitively, these $p$-values are based on the cumulative evidence from neighboring profiles $X_{j}\preceq X_{i}$, comparing the observed number of depressed cases to what would be expected if $\eta(X_{i})<\tau$. This ensures that even with dependent tests across profiles, each $p_{i}$ is valid.
    Formally, let $I(x)=\{\,j:X_{j}\preceq x\,\}$, $n(x)=|I(x)|$, and denote the concomitant responses by $Y_{(1)}(x),\dots,Y_{(n(x))}(x)$. Define $S_{k}(x)=\sum_{j=1}^{k}Y_{(j)}(x)$. Then an anytime-valid $p$-value for testing $\eta(x)<\tau$ is
    $p_{\tau}(x)=\min_{1\leq k\leq n(x)}\frac{\tau^{\,S_{k}(x)}(1-\tau)^{\,k-S_{k}(x)+1}}{B(1-\tau;\,k-S_{k}(x)+1,\,S_{k}(x)+1)},$
    where $B(z;a,b)$ is the incomplete Beta function. This construction ensures that $\mathbb{P}(p_{\tau}(x)\leq\alpha\mid X)\leq\alpha$ whenever $\eta(x)<\tau$ (see muller2025isotonic, Lemma 22).
  2. (2)
    Logical structure and multiple testing.
    The set of hypotheses $\{H_{0,i}\}$ is not independent. By monotonicity, if $X_{i}\preceq X_{j}$, then $H_{0,j}\subseteq H_{0,i}$. To encode these logical relations, we construct a directed acyclic graph (DAG): $G=(I,E)$, where $I=\{1,\dots,m\}$ indexes the hypotheses. We include a directed edge $i\to j$ whenever $X_{j}\preceq X_{i}$, so that edges always point from a stronger profile to a weaker one. From $G$, we derive a polyforest $F$ by retaining at most one parent for each node. This choice is not unique and can be made, for example, at random, by selecting the closest parent in $\ell_{\infty}$ norm, or alternatively using other distance measures such as the $\ell_{2}$ norm. This sparsification reduces complexity while preserving logical consistency.
    Let $R^{\text{ISS}}_{\alpha}(G,p)$ denote the rejection set from the DAG testing algorithm. The procedure begins by distributing the global $\alpha$-budget across the root nodes of $F$, with each root $i$ receiving share
[…]
    A root node $i$ is rejected if $p_{\tau,i}\leq\alpha_{i}$, in which case all its ancestors in $G$ are also rejected. It’s worth noting that here all its ancestors are defined with respect to the full DAG $G$, not only within the same tree of the polyforest. Any node that is logically stronger (i.e., dominates $i$ under the partial order) is simultaneously rejected. After each iteration, the rejected nodes are removed from $F$, and the unused $\alpha$ is reallocated to the remaining roots. The process continues until no further rejections are possible. (see formal pseudocode in Algorithm 1, muller2025isotonic)
    Intuitively, the algorithm passes significance level downward through the hierarchy, where roots receive budget first, and rejected nodes release unused $\alpha$ to their descendants. Ancestors are rejected whenever a descendant is, ensuring monotonicity. The final rejection set $R^{\text{ISS}}_{\alpha}(G,p)$ therefore respects both the partial order and the global FWER.
  3. (3)
    Upward closure.
    Denote the set of rejected nodes from the polyforest testing procedure as $R_{\alpha}^{\text{ISS}}(G,p)$. The estimated subgroup is then defined as $\hat{A}=\{\,x:x\succeq X_{i}\text{ for some }i\in R_{\alpha}^{\text{ISS}}(G,p)\,\}.$ By construction, $\hat{A}$ is the upward closure of $R_{\alpha}^{\text{ISS}}(G,p)$ with respect to the partial order $\preceq$. This closure step ensures that the selected subgroup respects the monotonicity structure: once a particular exposure profile $X_{i}$ is identified as high-risk, all more severe profiles $x\succeq X_{i}$ must also be included.
Under this construction, ISS guarantees that with probability at least $1-\alpha$, no profile with $\eta(x)<\tau$ is included in $\hat{A}$, controlling the family-wise error rate (FWER). In applied terms, ISS provides a conservative but rigorous way to delineate ACE exposure patterns linked to elevated depression risk. Concretely, this means that the subgroups we identify consist only of ACE combinations whose estimated probability of depression exceeds the chosen threshold, while still capturing as many truly high-risk profiles as possible. In this way, ISS leverages both valid statistical guarantees and the monotonicity structure to yield interpretable subgroups of adverse childhood experiences most strongly associated with depression.